### v2

@@ 4 Proof of Constant Competitive Ratio
Proof 4.11
Proof of Theorem 4.10: We construct $\text{{Optimal}}_{\text{align}}$ through an iterative grouping process consisting of two phases: first partitioning requests into temporal groups $\{\mathcal{Z}_{g}\}$, then further decomposing each group into execution batches $\{\mathcal{W}_{h}\}$. This process of transforming Optimal into $\textup{{Optimal}}_{\mathrm{align}}$ is demonstrated in Figure 4.
@@
Figure 4: From Optimal to $\textup{{Optimal}}_{\mathrm{align}}$
The construction begins with the initial group $\mathcal{Z}_{1}$, defined as follows:
Let $\mathcal{A}_{1}$ denote the set of all requests initiated at time 0. We define
$\delta_{1}=\max\{p_{i}+o_{i}\mid r_{i}\in\mathcal{A}_{1}\},$
which establishes the maximum processing duration among initial requests. This yields our first group:
$\mathcal{Z}_{1}=\{r_{i}\in\mathcal{I}\mid p_{i}+o_{i}\leq\delta_{1}\},$
containing all requests whose total duration falls within this initial temporal bound.
For subsequent groups ($g\geq 2$), we define $\delta_{g}$ recursively as
$\delta_{g}=\max\{p_{i}+o_{i}\mid r_{i}\in\mathcal{A}_{g}\}-\sum_{j=1}^{g-1}\delta_{j},$
where $\mathcal{A}_{g}$ consists of requests initiated but not yet completed by time $\sum_{j=1}^{g-1}\delta_{j}$. The $g$-th group $\mathcal{Z}_{g}$ is constructed as
$\mathcal{Z}_{g}=\bigg{\{}r_{i}\in\mathcal{I}\,\bigg{|}\,\sum_{j=1}^{g-1}\delta_{j}<p_{i}+o_{i}\leq\sum_{j=1}^{g}\delta_{j}\bigg{\}}.$
The transformation begins by delaying the execution of each group $\mathcal{Z}_{g}$ ($g\geq 2$) by $5\sum_{j=2}^{g-1}\delta_{j}$ time units while maintaining the internal timing relationships among all requests within the group. For $g\geq 2$, we define the adjusted cutting time as
$t_{g}=6\sum_{j=2}^{g-1}\delta_{j},$
effectively scaling the original temporal partitioning by a factor of 6.
Next, we decompose each group $\mathcal{Z}_{g}$ into smaller sub-batches $\{\mathcal{W}_{h}\}$ through the following partition:
$\mathcal{Z}_{g}=\mathcal{W}_{d_{g-1}+1}+\mathcal{W}_{d_{g-1}+2}+\cdots+\mathcal{W}_{d_{g}},$
where $d_{g}=\sum_{j=1}^{g}c_{j}$ counts the cumulative batches up to group $g$, with $c_{j}$ denoting the number of batches in $\mathcal{Z}_{j}$. The decomposition proceeds by first sorting all requests in $\mathcal{Z}_{g}$ by their original initiation times. Tnhen, for each subsequent batch $\mathcal{W}_{d_{g-1}+k}$, we select the maximal subset of remaining requests with the earliest initiation times while ensuring the total memory requirement does not exceed capacity $M$. This process repeats until all requests in $\mathcal{Z}_{g}$ are assigned to batches. This process guarantees that all batches except possibly the final one ($\mathcal{W}_{d_{g}}$) nearly saturate memory $M$ by Definition 4.4.
We now replace each request’s original response length in $\mathcal{W}_{h}$ with the batch’s average response length:
$\bar{o}_{h}=\frac{1}{n_{h}}\sum_{i=1}^{n_{h}}o_{h,i},$
where $n_{h}$ is the number of requests in batch $\mathcal{W}_{h}$. Let $\text{OPT}_{\text{transform}}$ denote the total end-to-end latency after this transformation.
By Definition 4.4, we derive a lower bound on the memory utilization efficiency $\eta_{h}$ for any non-final batch $\mathcal{W}_{h}$ (where $h\neq d_{g}$):
$\eta_{h}\geq\frac{\sum_{r_{i}\in\mathcal{W}_{h}}(2s_{i}+o_{i})}{2M}>\frac{M-\epsilon(M)}{2M}=\frac{1}{2}.$ (3)
@@
Given that the original processing window for $\mathcal{Z}_{g}$ satisfies $\Delta t_{g}<\delta_{g-1}+\delta_{g}$, inequality 3 implies that the main batches $\mathcal{W}_{d_{g-1}+1},\ldots,\mathcal{W}_{d_{g}-1}$ complete within $2(\delta_{g-1}+\delta_{g})$ time units. Furthermore, the final batch $\mathcal{W}_{d_{g}}$, with all response lengths bounded by $\delta_{g-1}+\delta_{g}$, completes within $\delta_{g-1}+\delta_{g}$ time units. Consequently, the complete batch sequence $\{\mathcal{W}_{d_{g-1}+k}\}_{k=1}^{c_{g}}$ is guaranteed to finish within the time interval $t_{g}-3\delta_{g-1},t_{g}+3\delta_{g}]$.
Therefore, given any $\mathcal{I}$, we establish a lower bound on Optimal’s total end-to-end latency through the following derivation:
$\text{TEL}(\text{{Optimal}}_{\text{transform}};\mathcal{I})$ $=\sum_{h=1}^{\infty}\left(\bar{o}_{h}\cdot\sum_{j=h+1}^{\infty}n_{j}\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$=\sum_{g=1}^{\infty}\sum_{h=d_{g-1}+1}^{d_{g}}\left(\bar{o}_{h}\cdot\left(\epsilon\left(\sum_{j=d_{g}+1}^{\infty}n_{j}\right)+\sum_{j=d_{g}+1}^{\infty}n_{j}\right)\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$\leq 6\cdot\left(\sum_{g=1}^{\infty}\left(\delta_{g}\cdot\sum_{j=d_{g}+1}^{\infty}n_{j}+\sum_{r_{l}\in\mathcal{Z}_{g}}(p_{i}-t_{g})\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}\right)$
$=6\cdot\text{TEL}(\text{{Optimal}};\mathcal{I}).$
@@
Proof 4.13
Proof of Theorem 4.12 The key insight is that minimizing total end-to-end latency for a fixed makespan requires maximizing throughput in the earliest possible time intervals. This relationship is captured precisely by our throughput metric $\Phi(\mathcal{X}_{k})=n_{k}/\bar{o}_{k}$, which measures requests completed per unit time in batch $\mathcal{X}_{k}$. Notably, this metric is exactly the reciprocal of our quality metric $F(\mathcal{X}_{k})$, establishing a direct connection between our optimization objective and scheduling efficiency.
@@
Furthermore, the memory utilization efficiency comparison reveals a significant advantage: in Algorithm $\text{{Sorted-F}}_{\text{align}}$, each group $\mathcal{Y}_{m}$ achieves memory utilization efficiency $\eta_{m}>1/2$ (Inequality 2), whereas in $\text{{Optimal}}_{\text{align}}$, the scaled allocation limits the efficiency to $\eta_{[t_{g},t_{g+1}]}\leq 1/6$. This efficiency gap directly translates to a strictly smaller makespan for $\text{{Sorted-F}}_{\text{align}}$ compared to $\text{{Optimal}}_{\text{transform}}$.
@@
Proof 4.14
Proof of Theorem 4.1: Given any $\mathcal{I}$, we combine the guarantees from Theorems 4.2, 4.5, 4.8, 4.10, and 4.12 to derive the following chain of inequalities for the total end-to-end latency:
@@
$\text{CR}(\text{{Sorted-F}})=\sup_{\mathcal{I}}\frac{\text{TEL}(\text{{Sorted-F}};\mathcal{I})}{\text{TEL}(\text{{Optimal}};\mathcal{I})}<48.$
@@
5 Approximation Methods for Sorted-F
As described in Section 3, the Phase 1 scheduling of Sorted-F requires solving the combinatorial optimization problem $\mathcal{X}^{*}=\arg\min_{\mathcal{X}\subseteq\mathcal{I}}F(\mathcal{X})$ subject to memory constraints $M(\mathcal{X},p_{i}+o_{i})\leq M$. As exhaustive search is infeasible for large request sets, we develop four approximation methods spanning the scalability spectrum. Each method makes distinct trade-offs between optimality guarantees and computational complexity.
@@ 5.1 Exact Dynamic Programming
The reverse traversal order is essential to prevent multiple inclusions of the same request while preserving the optimal substructure property. Path reconstruction occurs after all state transitions complete, converting stored indices into actual request objects. The $O(n^{2}M)$ complexity provides a practical exact solution for moderate-scale deployments where $n\leq 100$, establishing the fundamental benchmark for optimal scheduling decisions.
5.2 Scaled Dynamic Programming
This method provides a memory-efficient approximation to the exact combinatorial optimization problem. By quantizing memory usage values, it achieves polynomial complexity while maintaining $(1+\epsilon)$-optimality guarantees. The core innovation is a scaled representation of memory constraints:
$\hat{m}_{i}$ $=\left\lfloor\frac{s_{i}+o_{i}}{\lambda}\right\rfloor\quad\text{with}\quad\lambda=\frac{\epsilon M}{B}$
$\hat{M}$ $=\left\lfloor\frac{M}{\lambda}\right\rfloor,$
where $B$ controls precision. This reduces the state space dimension by factor $\lambda$, transforming the optimization to
$\mathcal{X}^{*}_{approx}=\arg\min_{\mathcal{X}\subseteq\mathcal{I}}$ $\frac{\sum_{r_{i}\in\mathcal{X}}o_{i}}{|\mathcal{X}|^{2}}$
subject to $\sum_{r_{i}\in\mathcal{X}}\hat{m}_{i}\leq\hat{M}$.
The complete pseudocode is provided in Algorithm 3 in Appendix 10.
The space-time complexity is $O(nB/\epsilon)$ versus $O(n^{2}M)$ for exact DP, enabling scaling to $n\leq 200$. The approximation quality exhibits a controllable trade-off: smaller $\epsilon$ values yield tighter bounds $F_{approx}\leq(1+\epsilon)F^{*}$ at the cost of $O(1/\epsilon)$ computation increase. This approach provides a theoretically grounded framework for balancing accuracy and efficiency in large-scale scenarios.
@@ 5.3 Local Swap Search
$\mathcal{X}_{k+1}=\begin{cases}\mathcal{X}_{k}&\nexists\text{ improving swap}\\[5.0pt] \mathcal{X}_{k}\oplus(r_{in},r_{out})&F(\mathcal{X}_{k}\oplus(r_{in},r_{out}))<F(\mathcal{X}_{k}),\end{cases}$
@@
The $O(n^{2})$ complexity enables deployment at scale $n\leq 500$. While lacking global optimality guarantees, the method demonstrates strong practical performance, yielding high-quality solutions that approach optimum scheduling efficiency with significantly reduced computational overhead compared to exact methods.
@@ 5.4 Quantile Greedy Selection
The $O(n)$ complexity ensures practical deployment for large-scale systems with $n\geq 1000$ requests. Phase 1 prioritizes predictable requests within distributional norms, while Phase 2 maximizes marginal utility of memory usage through the $o_{i}/(s_{i}+o_{i})$ efficiency metric. This combined approach maintains low computational overhead while effectively constraining outlier impact on scheduling performance.
@@ 5.5 Algorithm Selection Guidelines
The four approximation algorithms present distinct operational characteristics that enable optimized deployment across the scalability spectrum. System designers should consider the algorithmic trade-offs in terms of solution quality, computational efficiency, and applicability range when selecting schedulers for specific deployment scenarios. The comparative properties are formally summarized in Table 3, which serves as the primary reference for implementation decisions.
Table 3: Algorithmic Properties and Operational Ranges
Algorithm Optimality Complexity Recommended Scale
Exact Dynamic Programming Optimal $O(n^{2}M)$ $n\leq 100$
Scaled DP $(1+\epsilon)$-approximation $O(nB/\epsilon)$ $n\leq 200$
Local Swap Search Local optimum $O(n^{2})$ $n\leq 500$
@@
Selection decisions should prioritize alignment between algorithm capabilities and deployment requirements, with accuracy-sensitive systems favoring exact methods despite computational costs, while latency-critical large-scale deployments benefit from linear-time heuristics. For moderate-scale systems with $200\leq n\leq 500$, the local search approach offers a balanced compromise, delivering near-optimal solutions with quadratic complexity that remains computationally manageable. Environmental factors including workload heterogeneity and available hardware resources further influence implementation choices; heterogeneous systems particularly benefit from the Quantile Greedy’s statistical sampling approach that constrains outlier impact through distribution-aware selection. When considering temporal performance profiles, designers should note that while worst-case complexity bounds determine theoretical scaling, practical implementations frequently achieve substantially better performance, particularly for the Local Swap method which empirically converges in constant iterations across diverse workloads. Hybrid implementations can dynamically switch algorithms based on real-time request volume to maintain optimal scheduling efficiency across varying load conditions. In the next section, we will compare the performance of these approximation methods on real-data numerical experiments.

### v3

@@ 4 Proof of Constant Competitive Ratio
Definition 4.16 (Time-window scaling and response-length aligning $\textup{{Optimal}}\mapsto\textup{{Optimal}}_{\mathrm{align}}$)
Let Optimal have initiation times $\{p_{i}\}$ and response lengths $\{o_{i}\}$. We construct $\textup{{Optimal}}_{\mathrm{align}}$ in three steps.
(i) Temporal grouping. Define $\mathcal{A}_{1}:=\{r_{i}:\,p_{i}=0\}$ and
$\delta_{1}:=\max\{p_{i}+o_{i}\mid r_{i}\in\mathcal{A}_{1}\},\qquad\mathcal{Z}_{1}:=\{r_{i}\in\mathcal{I}\mid p_{i}+o_{i}\leq\delta_{1}\}.$
For $g\geq 2$, let $S_{g-1}:=\sum_{j=1}^{g-1}\delta_{j}$ and let $\mathcal{A}_{g}$ be the set of requests initiated but not yet completed by time $S_{g-1}$. Define
$\delta_{g}:=\max\{p_{i}+o_{i}\mid r_{i}\in\mathcal{A}_{g}\}-S_{g-1},\qquad\mathcal{Z}_{g}:=\{r_{i}\in\mathcal{I}\mid S_{g-1}<p_{i}+o_{i}\leq S_{g-1}+\delta_{g}\}.$
(ii) Time-window scaling. For each $g\geq 2$, shift all requests in $\mathcal{Z}_{g}$ forward by $5\cdot\sum_{j=2}^{g-1}\delta_{j}$ time units, preserving the within-group relative timing. Define the (scaled) cutting time
$t_{g}:=6\cdot\sum_{j=2}^{g-1}\delta_{j},$
with the convention that an empty sum equals 0.
(iii) Batching + response-length aligning. Within each (shifted) group $\mathcal{Z}_{g}$, sort requests by their (shifted) initiation times and greedily partition them into batches
$\mathcal{Z}_{g}=\mathcal{W}_{d_{g-1}+1}+\cdots+\mathcal{W}_{d_{g}},$
where each batch $\mathcal{W}_{h}$ is the maximal prefix (in this order) whose total memory requirement does not exceed $M$. Hence, all batches except possibly the final one in each group nearly saturate memory $M$ by Definition 4.6. Finally, for each batch $\mathcal{W}_{h}$ with $n_{h}$ requests and response lengths $\{o_{h,i}\}_{i=1}^{n_{h}}$, replace all response lengths in the batch by the average
$\bar{o}_{h}:=\frac{1}{n_{h}}\sum_{i=1}^{n_{h}}o_{h,i}.$
We denote the resulting analytically transformed schedule by $\textup{{Optimal}}_{\mathrm{align}}$.
Figure 5 schematically illustrates Definition 4.16.
@@
Figure 5: Schematic illustration of Definition 4.16 (from Optimal to $\textup{{Optimal}}_{\mathrm{align}}$).
Proof 4.17
Proof of Theorem 4.15: We now prove $\mathrm{TEL}(\textup{{Optimal}}_{\mathrm{align}};\mathcal{I})\leq 6\cdot\mathrm{TEL}(\textup{{Optimal}};\mathcal{I})$.
For group $\mathcal{Z}_{g}$, let $\Delta t_{g}$ denote its original execution window length under Optimal:
$\Delta t_{g}:=\max_{r_{i}\in\mathcal{Z}_{g}}(p_{i}+o_{i})-\min_{r_{i}\in\mathcal{Z}_{g}}p_{i}.$
By construction of $\{\mathcal{Z}_{g}\}$ (Definition 4.16(i)), all requests in $\mathcal{Z}_{g}$ complete no later than $S_{g-1}+\delta_{g}$, while $\mathcal{A}_{g}$ contains requests alive at time $S_{g-1}$; in particular, $\max_{r_{i}\in\mathcal{A}_{g}}(p_{i}+o_{i})=S_{g-1}+\delta_{g}$. This implies the crude bound $\Delta t_{g}<\delta_{g-1}+\delta_{g}$, which is the only property we will use.
By Definition 4.6, every non-final batch $\mathcal{W}_{h}$ (i.e., $h\neq d_{g}$ within its group) satisfies
$\eta_{h}\;\geq\;\frac{\sum_{r_{i}\in\mathcal{W}_{h}}(2s_{i}+o_{i})}{2M}\;>\;\frac{M-L(M)}{2M}\;\to\;\frac{1}{2}.$ (6)
@@
Using (6) and the window bound $\Delta t_{g}<\delta_{g-1}+\delta_{g}$, the main batches $\mathcal{W}_{d_{g-1}+1},\ldots,\mathcal{W}_{d_{g}-1}$ complete within $2(\delta_{g-1}+\delta_{g})$ time units. The final batch $\mathcal{W}_{d_{g}}$ has all (aligned) response lengths bounded by $\delta_{g-1}+\delta_{g}$, hence completes within $\delta_{g-1}+\delta_{g}$ time units. Therefore, the whole batch sequence of group $g$ finishes within $t_{g}-3\delta_{g-1},\;t_{g}+3\delta_{g}]$, where $t_{g}$ is the scaled cutting time in Definition 4.16(ii).
It remains to convert the above window bound into a TEL bound. We rewrite $\mathrm{TEL}(\texttt{Optimal}_{\mathrm{align}};\mathcal{I})$ in batch form, regroup by $\{\mathcal{Z}_{g}\}$, and use that group $g$ completes within $t_{g}-3\delta_{g-1},\,t_{g}+3\delta_{g}]$. The next lemma formalizes this bookkeeping step.
Lemma 4.18
Given the transformed schedule $\textup{{Optimal}}_{\mathrm{align}}$ constructed in Definition 4.16, we have
@@
Proof 4.20
Proof of Theorem 4.19: The key insight is that minimizing total end-to-end latency for a fixed makespan requires maximizing throughput in the earliest possible time intervals. This relationship is captured precisely by our throughput metric $\Phi(\mathcal{X}_{k})=n_{k}/\bar{o}_{k}$, which measures requests completed per unit time in batch $\mathcal{X}_{k}$. Notably, this metric is exactly the reciprocal of our quality metric $F(\mathcal{X}_{k})$, establishing a direct connection between our optimization objective and scheduling efficiency.
@@
Furthermore, the memory utilization efficiency comparison reveals a significant advantage: in Algorithm $\text{{Sorted-F}}_{\text{align}}$, each group $\mathcal{Y}_{m}$ achieves $\eta_{m}\geq 1/2$ by (5), whereas in $\text{{Optimal}}_{\text{align}}$, the scaled allocation limits the efficiency to $\eta_{[t_{g},t_{g+1}]}\leq 1/6$. This efficiency gap directly translates to a strictly smaller makespan for $\text{{Sorted-F}}_{\text{align}}$ compared to $\text{{Optimal}}_{\text{align}}$.
@@
Proof 4.21
Proof of Theorem 4.1: Given any $\mathcal{I}$, we combine the guarantees from Theorems 4.4, 4.7, 4.13, 4.15, and 4.19 to derive the following chain of inequalities for the total end-to-end latency:
@@
$\text{CR}(\text{{Sorted-F}})<48.$
@@
5 Approximation Algorithm and Practical Implementation under Bad Predictions
As described in Section 3, Phase 1 of Sorted-F requires solving the combinatorial optimization problem $\mathcal{X}^{*}=\arg\min_{\mathcal{X}\subseteq\mathcal{I}}F(\mathcal{X})$ subject to the memory feasibility constraints. When this Phase 1 subproblem is solved exactly, the resulting schedule corresponds to the idealized Sorted-F analyzed in Section 4. However, exhaustive search is infeasible for large request sets, so we present one exact solver, such as dynamic programming, as a benchmark, and two scalable heuristics for practical use. The heuristic methods are included for empirical performance and runtime considerations; we do not claim worst-case approximation guarantees for them, nor do we claim that the constant-factor bound in Section 4 carries over when Phase 1 is implemented heuristically.
@@ 5.1 Exact Dynamic Programming
The reverse traversal order is essential to prevent multiple inclusions of the same request while preserving the optimal substructure property. Path reconstruction occurs after all state transitions complete, converting stored indices into actual request objects. The $O(n^{2}M)$ complexity provides a practical exact solution for moderate-scale deployments where $n\leq 100$, establishing the fundamental benchmark for Phase 1 batch selection. In particular, when Phase 1 is solved exactly, the resulting Sorted-F schedule matches the algorithmic assumption underlying the constant-factor analysis in Section 4.
@@ 5.2 Local Swap Search
$\mathcal{X}_{k+1}=\begin{cases}\mathcal{X}_{k}&\nexists\text{ improving swap}\\ \mathcal{X}_{k}\oplus(r_{in},r_{out})&F(\mathcal{X}_{k}\oplus(r_{in},r_{out}))<F(\mathcal{X}_{k}),\end{cases}$
@@
The $O(n^{2})$ complexity enables deployment at scale $n\leq 500$. This method is a heuristic: it may get trapped in local minima and we do not provide a worst-case approximation guarantee for the Phase 1 objective. Consequently, when Phase 1 is implemented via Local Swap Search, we do not claim that the constant competitive-ratio bound proved in Section 4 continues to hold.
@@ 5.3 Quantile Greedy Selection
The $O(n)$ complexity ensures practical deployment for large-scale systems with $n\geq 1000$ requests. This method is designed as a lightweight heuristic based on distributional trimming and greedy filling; we do not provide a worst-case approximation guarantee for the Phase 1 objective. Accordingly, when Phase 1 is implemented via Quantile Greedy Selection, we do not claim that the constant competitive-ratio bound proved in Section 4 continues to hold.
@@ 5.4 Algorithm Selection Guidelines
The three approximation algorithms present distinct operational characteristics that enable optimized deployment across the scalability spectrum. System designers should consider the algorithmic trade-offs in terms of solution quality, computational efficiency, and applicability range when selecting schedulers for specific deployment scenarios. The comparative properties are formally summarized in Table 3, which serves as the primary reference for implementation decisions.
Table 3: Algorithmic Properties and Operational Ranges (Phase 1 subproblem solvers)
Algorithm Guarantee Complexity Recommended Scale
Exact Dynamic Programming Exact optimum $O(n^{2}M)$ $n\leq 100$
Local Swap Search Heuristic $O(n^{2})$ $n\leq 500$
@@
Selection decisions should prioritize alignment between algorithm capabilities and deployment requirements. Accuracy-sensitive systems may favor the exact DP despite its computational cost, while latency-critical large-scale deployments benefit from linear-time heuristics. For moderate-scale systems with $200\leq n\leq 500$, the local search approach offers a balanced compromise with quadratic complexity that remains computationally manageable. Environmental factors including workload heterogeneity and available hardware resources further influence implementation choices; heterogeneous systems may particularly benefit from the Quantile Greedy’s distribution-aware selection that helps constrain outlier impact. We emphasize that the Local Swap and Quantile Greedy procedures are included primarily for computational efficiency and observed performance; we do not provide worst-case approximation guarantees for them, nor do we claim that the constant-factor bound in Section 4 applies when Phase 1 is solved heuristically. Hybrid implementations can dynamically switch algorithms based on real-time request volume to maintain robust performance across varying load conditions. In the next section, we will compare the performance of these methods on real-data numerical experiments.
5.5 Handling Prediction Uncertainty via Adaptive Output Length Refinement
The algorithms presented thus far assume that output lengths $o_{i}$ are known exactly. In practice, output lengths must be predicted, and predictions are inherently imperfect. We now describe how to extend Sorted-F to operate robustly under prediction uncertainty by incorporating the adaptive refinement strategy of Chen et al. (2025).
Prediction Model.
For each request $i$, a machine learning predictor provides an interval $\ell_{i},u_{i}]$ such that the true output length satisfies $o_{i}\in[\ell_{i},u_{i}]$. Such interval predictions arise naturally from classification-based predictors that assign requests to output-length bins. The ratio $\alpha:=\inf_{i}\ell_{i}/u_{i}\in(0,1]$ captures prediction confidence; smaller $\alpha$ indicates greater uncertainty.
Challenges with Naïve Approaches.
Two straightforward strategies are to substitute $o_{i}$ with either $u_{i}$ (conservative) or $\ell_{i}$ (optimistic) throughout the algorithm. The conservative approach, termed $\mathcal{A}_{\max}$ in Chen et al. (2025), prevents memory overflow but suffers from severe performance degradation as prediction accuracy decreases—memory is over-reserved, reducing throughput. The optimistic approach risks memory violations when actual output lengths exceed predictions.
Adaptive Refinement via $\mathcal{A}_{\min}$.
The key insight of Chen et al. (2025) is to start optimistically and adapt during inference. Their algorithm $\mathcal{A}_{\min}$ initially treats each request’s output length as the lower bound $\ell_{i}$. During inference, if request $i$ has generated $\tilde{o}_{i}$ tokens without completing, its effective output length estimate is updated to $\tilde{o}_{i}$. This dynamic refinement allows the scheduler to react to actual generation progress rather than committing to potentially inaccurate predictions. When memory overflow is projected, $\mathcal{A}_{\min}$ employs ordered eviction: requests with smaller current estimates $\tilde{o}_{i}$ are removed from the active batch first, as they are closer to completion and their eviction frees memory with minimal wasted computation.
Integration with Sorted-F.
We integrate adaptive refinement into Sorted-F as follows. In Phase 1, batch construction uses the predicted lower bounds $\ell_{i}$ in place of $o_{i}$ when evaluating the objective $F(\mathcal{X})$ and checking memory feasibility. In Phase 2, the scheduling execution maintains a running estimate $\tilde{o}_{i}$ for each active request, initialized to $\ell_{i}$ and updated to $\max\{\tilde{o}_{i},t-p_{i}\}$ at each time step $t$, where $p_{i}$ is the start time of request $i$. Memory feasibility checks use these running estimates. If at any time step the projected memory usage exceeds $M$, ordered eviction is triggered: requests are removed from the active set in increasing order of $\tilde{o}_{i}$ until feasibility is restored. Evicted requests are returned to the pending queue and may be rescheduled in subsequent batches.
[…]